# Grid-based Multi-Robot Coverage Path Planning

### v1

@@ 2 $\mathbf{x}_{\text{last}}\leftarrow\left(v_{1},0\right)$,  $j=2$,  $b\leftarrow 1$
3 $P\leftarrow\left(\mathbf{x}_{\text{last}},\right)$,  $S_{\text{goal}}\leftarrow(\mathbf{x}^{(\pi)}_{j}=(v_{j},\cdot),\,)$
@@
5       find a plan from $\mathbf{x}_{\text{last}}$ to $S_{\text{goal}}$ via ML-SIPP
6       if a valid plan $\tilde{P}$ is found then
7             append $\tilde{P}$ to the end of $P$
@@ 8             $j\leftarrow j+1$,  $b\leftarrow 1$
9             $S_{\text{goal}}\leftarrow(\mathbf{x}^{(\pi)}_{j}=(v_{j},\cdot),\,)$
@@
11       else if $b\leq b_{\text{max}}$ or $P$ is not empty then
12             pop the last state from $P$ until $\mathbf{x}_{j-b}^{(\pi)}$
13             $S_{\text{goal}}\leftarrow(\mathbf{x}_{j-b}^{(\pi)},\,)+S_{\text{goal}}$,  $b\leftarrow b+1$
@@
16             $\mathbf{x}_{\text{last}}\leftarrow(v_{1},0)$, $S_{\text{goal}}\leftarrow(\mathbf{x}_{1}^{(\pi)},...,\mathbf{x}_{|\pi|}^{(\pi)})$
17             return plan from $\mathbf{x}_{\text{last}}$ to $S_{\text{goal}}$ via ML-SIPP
@@
20      $\mathbf{x}_{\text{last}}\leftarrow$ the last state in $P$
@@ VII Experiments and Results
This section describes our numerical results of simulations on an Ubuntu 22.04 PC with a 2.50GHz Intel® Core® i5-13490F CPU and 32GB RAM. Our proposed algorithms and the baseline methods are publicly available at Github [^1] https://github.com/reso1/LS-MCPP, with all the benchmark instances and reported numerical results provided as well. In Sec. VII-E, we deployed the proposed algorithms on two Pepper [^2] https://www.aldebaran.com/en/pepper robots in an indoor environment to validate our planning pipeline for MCPP.
@@ VII-A Experiment Setup
Instances: We generate a set of instances to benchmark different methods utilizing the four-way connected 2d undirected grid graphs from previous grid-based MCPP works [tang2023mixed, tang2024large] as well as the 2D pathfinding benchmark [sturtevant2012benchmarks]. As shown in Fig. 13, the graphs in row 1 have unweighted edges with a uniform weight of 1, whereas the graphs in row 5 have weighted edges whose weights are randomly generated from 1 to 3. Originally, each graph can be contracted to a hypergraph where all its hypervertices are complete. The original graphs are then used to generate 12 groups of 12 mutated MCPP instances (i.e., mutations) of $k$ robots, resulting from a removal ratio $\rho=0,1,...,11$ and a seed $s=0,1,...,11$. Each mutation $(G,I,R)$ with $\rho$ and $s$ is constructed by (1) randomly sampling (with seed $s$) a set of $k$ vertices from $G$ as $R$ and (2) randomly removing (with seed $s$) vertices from $G$ by $x=(\frac{100}{12}\cdot\rho)\%$, meanwhile ensuring the root vertex of any robot and all its neighbors are not removed. The root vertices in $R$ are sampled to ensure they are not cut vertices for constructing well-formed instances [vcap2015complete, ma2019searching], that is, a robot can wait indefinitely at its root vertex without blocking any other robot. Note that the graph edge weights of each mutation are also randomly generated using its corresponding seed $s$. For all mutations, we set $C=0.5$ as the $90^{\circ}$ turning cost as described in Sec. LABEL:sec:turn_cost.
@@
Baselines: We describe all the methods being compared in our experiments. (1) Single-Robot CPP: To verify the effectiveness of our ESTC with the proposed hyperedge weight definition, we use a baseline method, namely ESTC-UW, that computes the circumnavigating coverage path on a spanning tree assuming the edges are unweighted. We also performed an ablation study on the two local optimizations for ESTC: parallel rewiring (PR) in Sec. IV-A and turn reduction (TR) in Sec. IV-C. (2) MCPP: As mentioned in Sec. IV-D, the proposed ESTC algorithm can be integrated into various grid-based MCPP methods, making them immediately applicable to MCPP. To reiterate, VOR, MFC, MIP, and MSTC are the comparing baseline methods with our LS-MCPP. In addition, we also compare LS-MCPP with and without the vertex-wise operators that are new to our previous work [tang2024large]  [^3] In the ablation study of our previous work [tang2024large], we have shown the effectiveness of grow, deduplicate, and exchange operators as well as the use of Forced Deduplication (Alg. 3) and heuristic sampling for LS-MCPP., namely LS(+VO) and LS(-VO) respectively. Note that all the methods compared use ESTC with TR and PR applied by default. (3) Deconflicted MCPP: Given LS(+VO) solutions to produce the initial time-embedded trajectories, we compare the deconflicting performance with the chaining approach (CHA), the multi-label approach (MLA), and the adaptive approach (ADA) as the low-level planner and PBS as the high-level planner.
@@ VII-B Ablation Study
Figure 13: Ablation Study on 6 unweighted graphs (row 1) and 6 weighted graphs (row 5). Each plot (rows 2-4 and 6-8) includes 12 groups of mutated instances with $k$ robots, each contains 12 mutations that randomly remove a portion (x-axis) of vertices, with random root vertices and edge weights (see Sec. VII-A). Single-Robot CPP (rows 2 and 6): mean and standard deviation (std) of cost reductions of ESTC variants over ESTC-UW. MCPP (rows 3 and 7): mean and std of makespan reductions of MCPP methods over VOR. Deconflicted MCPP (rows 4 and 8): mean and std of makespan increments over the input LS-MCPP solution for mutations where all approaches have at least one mutation succeeded (left y-axis) and their success rates in all 12 mutations (right y-axis). Note that when one or two approaches fail at all 12 mutations, we also plot for the mutations where the other approach(es) all succeed.
@@
Single-Robot CPP: Comparing all four ESTC variants with the baseline ESTC-UW, we observe a more critical cost reduction on weighted graphs (row 6 of Fig. 13) than on unweighted graphs (row 2 of Fig. 13). For unweighted graphs, TR results in significant performance improvements since uniformly weighted edges can be aligned in one direction to minimize turning costs. For weighted graphs, TR is much less effective as there are few tie-breaking cases for reducing the turning costs, while PR achieves higher cost reductions as the vertex removal percentage increases. In general, ESTC shows its effectiveness for single-robot CPP, and both PR and TR contribute to substantial path cost reductions.
@@
Figure 14: LS-MCPP solutions with increasing max iterations $M=s\times\sqrt{|V|/k}$. Colored lines represent different values of $s$.
@@
MCPP: We compare the solution quality of MCPP between different methods in rows 3 and 7 of Fig. 13. Generally speaking, the mutated instances become more complex as the vertex removal ratio increases. Our LS-MCPP consistently outperforms other methods across all instances, while the performance of the other methods varies, with each showing advantages in different instances. Note that due to the huge memory and runtime requirements of MIP, we only test it on relatively smaller instances (i.e., names with a suffix of small or medium), and its performance is highly restricted by the runtime limit of 1 hour. Specifically for our LS-MCPP, as the vertex removal ratio increases, LS(+VO) performs better with the additional vertex operators compared to LS(-VO). However, when the instances become more complex with a large map and more robots, the solution quality of LS(+VO) gets closer to LS(-VO) due to an inadequate number of local search iterations limited by $M$. To demonstrate that we can further improve the solution quality by saturating LS(+VO) with more iterations, we compare the LS(+VO) solutions with increasing maximum numbers of iterations, as shown in Fig. 14. The solution quality steadily improves as the maximum number $M$ of iterations increases with a linear increase in runtime. In practice, $M$ can be flexibly adjusted according to the available runtime budget, allowing a trade-off between better solution quality and faster planning.
Figure 15: Deconflicting different MCPP solutions using PBS as the high-level planner and our adaptive approach as the low-level planner.
Deconflicted MCPP: We explain the Deconflicted MCPP ablation study results in Fig. 13 (rows 4 and 8). CHA demonstrates the highest makespan increments and the lowest success rate due to its inability to backtrack to previous goal states, resulting in incompleteness. This limitation makes CHA susceptible to ”dead-ends” when no valid succeeding states are available. Consequently, it reports a higher number of failure nodes during PBS search, which either prevents PBS from finding a valid priority ordering or significantly increases the number of PBS nodes explored (sometimes exceeding the runtime limit in large-scale instances). MLA produces high-quality trajectory sets, with no makespan increases for most of the ”easier” instances it can handle within the runtime limit. However, for more complex instances involving larger maps and more robots, the added label dimension across the entire coverage path expands the search space considerably, which significantly increases planning time for each PBS node and results in a lower success rate due to timeouts. Our ADA succeeds in all instances, achieving high solution quality with no makespan increases in most cases. This is due to its ability to restrict the label-augmented search space within a local window, rather than across the entire coverage path as MLA, and only when backtracking to previous states is necessary. In Fig. 15, we also compare the deconflicting results of different MCPP solutions for floor_large and terrain_large using PBS and ADA, following the same plotting conventions as in rows 4 and 8 of Fig. 13. The results demonstrate the better solution quality of the proposed LS-MCPP when used as the input MCPP solution for deconfliction. In contrast, as MFC and MSTC* have larger makespans due to the higher overlapping covered regions across the map, the deconflicting procedures on them are more time-consuming. Notably, there is one failure case of deconflicting the MFC solution where the one-hour runtime limit is exceeded. For VOR, as there are no overlapping regions in its solution, PBS immediately returns the initial time-embedded trajectories without any conflict.
VII-C Planning Runtime Comparison
We report the planning runtime of all the MCPP methods and the Deconflicted MCPP methods described in Sec. VII-A across all the 12 groups of instances mutated from the grid graphs shown in rows 1 and 5 of Fig. 13. We show the planning runtime comparison of methods for MCPP in Tab. I, where our LS-MCPP requires more time to explore the solution neighborhood. In particular, adding vertex-wise operators only increases the runtime of LS-MCPP by a little if we compare LS(-VO) and LS(+VO). Nevertheless, it produces high-quality MCPP solutions within around 13 minutes even for large-scale instances such as NewYork1 and Shanghai2, which is acceptable for offline planning in real-world robotics applications.
@@ VII-D Physical Robot Experiments
Figure 16: Demonstration of Cayenne and Cumin performing the coverage task cooperatively without conflict. Four frames from the task execution are illustrated with their trajectories and the graph shown on the left side. Static obstacles, trajectory of Cayenne, and trajectory of Cumin are highlighted in red, blue, and yellow respectively. The entire planning pipeline was completed in 0.1 seconds and the coverage task was completed in approximately 6 minutes.
@@
We constructed an MCPP instance of an indoor environment with a Vicon [^4] https://www.vicon.com motion capture system for localization as shown in Fig. 17. Our pipeline first computes the conflict-free set of trajectories for the robots by first finding an MCPP solution via LS-MCPP and then deconflicting the MCPP solution via PBS and ADA. After offline planning is completed, the trajectories are sent to the robots to perform the coverage task through a centralized scheduler written in ROS [^5] https://www.ros.org. Fig. 17 demonstrates the execution procedure of the resulting trajectories using our planning pipeline. The detailed setup is described as follows.
@@
Multi-Robot System: We employ two Pepper robots, named Cayenne and Cumin, each with a square bounding box of 0.55 meters. The 2D grid graph represents the indoor environment as a $10\times 6$ map with each grid unit sized at 0.65 meters to accommodate their bounding boxes. The initial root vertices of Cayenne and Cumin are placed at the bottom-left and top-right grids of the graph, respectively. Both Peppers can have omnidirectional movement driven by velocity commands $u=(v_{x},v_{y},\omega)$. The linear velocities $v_{x}$ and $v_{y}$ are capped at 0.15 meters per second, while the angular velocity $\omega$ is restricted to 0.5 radians per second. To determine Pepper’s non-uniform traversal costs in the graph (i.e., the edge weights), we recorded the maximum and minimum time costs to traverse a single grid unit by performing random move actions across the graph. We then sampled weights randomly within this range and assigned them to the edges. Although Peppers can move omnidirectionally, we assume they always face the moving directions by integrating turn actions as described in Sec. LABEL:sec:turn_cost. We set $C$ as the maximum action time of 90 rotations for Peppers. In the final constructed graph, the time costs for move actions range from 7.2 to 10.8 seconds, while the time cost for a 90 turn is 6.0 seconds.
@@
Centralized Scheduler: Given the set of conflict-free trajectories resulted from our planning pipeline, we generate an action sequence consisting of move, turn, wait actions between consecutive states for each Pepper. A centralized scheduler is employed to sequentially send actions to the corresponding Pepper, maintaining a positional tolerance of 0.01 meters and a rotational tolerance of 0.02 radians. The velocity commands for each action are generated using a proportional controller operating at 20 Hz to align with the nominal action time cost. In practice, the move and turn action time costs are not precise for the Peppers, as we assume instantaneous velocity changes and do not account for their real-time dynamics or potential perturbations. This often leads to an overestimation of the actual action time costs, as the time for acceleration and deceleration is ignored, though underestimations can also occur occasionally. Therefore, we make the centralized scheduler to synchronize the discrepancies between the nominal time costs $t$ and the actual time costs $t^{\prime}$ for the actions. Upon completing each action, Pepper sends the time difference $\Delta t=t-t^{\prime}$ back to the scheduler. If $\Delta t<0$, the scheduler signals Pepper to wait for a period of $-\Delta t$ to match the nominal time cost; otherwise, the scheduler instructs the other Pepper to add $\Delta t$ to its waiting period after completing its current action.
@@
Our centralized scheduler currently uses a basic approach to synchronize time differences between nominal and actual action time costs, which may lead to increased coverage task completion times. More advanced techniques from MAPF execution research, such as constructing an action dependency graph [honig2019persistent], could address this issue. However, exploring these advanced methods is beyond the scope of this paper.
VII-E Physical Robot Experiments
Figure 17: Demonstration of Cayenne and Cumin performing the coverage task cooperatively without conflict. Four frames from the task execution are illustrated with their trajectories and the graph shown on the left side. Static obstacles, trajectory of Cayenne, and trajectory of Cumin are highlighted in red, blue, and yellow respectively. The entire planning pipeline was completed in 0.1 seconds and the coverage task was completed in approximately 6 minutes.
We deployed the proposed algorithmic pipeline for MCPP on two Pepper [^6] https://www.aldebaran.com/en/pepper robots in an indoor environment with a Vicon [^7] https://www.vicon.com/ motion capture system for localization. Given the grid graph of the indoor environment, the pipeline computes the conflict-free set of plans for the robots by first finding an MCPP solution via LS-MCPP and then deconflicting the MCPP solution via PBS and ADA. After offline planning is completed, the plans are sent to the robots to perform the coverage task through a centralized scheduler written in ROS [^8] https://www.ros.org/. Fig. 17 demonstrates the execution procedure of the resulting plans using our planning pipeline, and the detailed experiment setup is described as follows.
Multi-Robot System: We employ two Pepper robots, named Cayenne and Cumin, each with a square bounding box of 0.55 meters. The 2D grid graph represents the indoor environment as a $10\times 6$ map with each grid unit sized at 0.65 meters to accommodate their bounding boxes. The initial root vertices of Cayenne and Cumin are placed at the bottom-left and top-right grids of the graph, respectively. Both Peppers are capable of omnidirectional movement (i.e., holonomic) driven by velocity commands $u=(v_{x},v_{y},\omega)$. The linear velocities $v_{x}$ and $v_{y}$ are capped at 0.15 meters per second, while the angular velocity $\omega$ is restricted to 0.5 radians per second. To determine Pepper’s non-uniform move action costs in the graph (i.e., the edge weights $w_{e}$), we recorded the maximum and minimum time costs to traverse a single grid unit by performing random move actions across the graph. We then sampled weights randomly within this range and assigned them to the edges. We set $C$ as the maximum time required for 90 rotations. In the final constructed graph, the time costs for move actions range from 7.2 to 10.8 seconds, while the time cost for a 90 turn is 6.0 seconds.
Centralized Scheduler: Given the set $\mathcal{P}$ of conflict-free plans resulted from our planning pipeline, we generate an action sequence consisting of move, turn, wait actions between consecutive states for each Pepper. A centralized scheduler is employed to sequentially send actions to the corresponding Pepper, maintaining a positional tolerance of 0.01 meters and a rotational tolerance of 0.02 radians. The velocity commands for each action are generated using a proportional controller operating at 20 Hz to align with the nominal action time cost. In practice, the move and turn action time costs are not precise for the Peppers, as we assume instantaneous velocity changes and do not account for their real-time dynamics or potential perturbations. This often leads to an overestimation of the actual action time costs, as the time for acceleration and deceleration is ignored, though underestimations can also occur occasionally. Therefore, we make the centralized scheduler to synchronize the discrepancies between the nominal time costs $t$ and the actual time costs $t^{\prime}$ for the actions. Upon completing each action, Pepper sends the time difference $\Delta t=t-t^{\prime}$ back to the scheduler. If $\Delta t<0$, the scheduler signals Pepper to wait for a period of $-\Delta t$ to match the nominal time cost; otherwise, the scheduler instructs the other Pepper to add $\Delta t$ to its waiting period after completing its current action.
Remark 6.
Our centralized scheduler currently uses a basic approach to synchronize time differences between nominal and actual action time costs, which may lead to increased coverage task completion times. More advanced techniques from MAPF execution research, such as constructing an action dependency graph, could address this issue. However, exploring these advanced methods is beyond the scope of this paper.

### v2

@@ 2 $\mathbf{x}_{\text{last}}\leftarrow\left(v_{1},0\right)$,  $j=2$,  $b\leftarrow 1$
3 $\tau\leftarrow\left(\mathbf{x}_{\text{last}},\right)$,  $S_{\text{goal}}\leftarrow(v_{j},\,)$
@@
5       find a trajectory from $\mathbf{x}_{\text{last}}$ to $S_{\text{goal}}$ via ML-SIPP
6       if a valid trajectory $\tilde{\tau}$ is found then
7             append $\tilde{\tau}$ to the end of $\tau$
@@ 8             $j\leftarrow j+1$,  $b\leftarrow 1$
9             $S_{\text{goal}}\leftarrow(v_{j},\,)$
@@
11       else if $b\leq b_{\text{max}}$ or $\tau$ is not empty then
12             pop the last state from $\tau$ until $v_{j-b}$
13             $S_{\text{goal}}\leftarrow(v_{j-b},\,)+S_{\text{goal}}$,  $b\leftarrow b+1$
@@
16             mark $N$ as postponed for multi-label approach
17             return $\emptyset$
@@
20      $\mathbf{x}_{\text{last}}\leftarrow$ the last state in $\tau$
@@
VI-D Integrating Turning Costs
So far, the trajectory computation assumes the robots are holonomic, that is, they can move omnidirectional without the need to change their headings between consecutive move actions. We now show that one can remove this assumption for non-holonomic robots by adding an orientation $\psi\in\{\text{E},\text{N},\text{W},\text{S}\}$ to the state space, restricted to East (E), North (N), West (W), and South (S) in the four-neighbor grid graph $G$. Correspondingly, we allow a robot to turn at vertex $v$ from $\psi$ to $\psi^{\prime}\neq\psi$ at a time cost of $t^{\prime}-t=C\cdot\frac{|\psi^{\prime}-\psi|}{90^{\circ}}>0$, where $C$ is a constant cost for a $90^{\circ}$ turns. Effectively, each robot performs a turn-wait-move action. With this modification, one can deconflict an MCPP solution for non-holonomic robots.
Additionally, the above modification requires the parallel rewiring optimization for ESTC to consider the turn costs when calculating the path cost difference (Sec. IV-C). Other components of LS-MCPP (Alg. 2), including ESTC (Alg. 1) and its turn reduction optimization, all remain unchanged.
Figure 11: MCPP instances with $k$ robots for the simulation experiments. Top row: unweighted grid graphs. Bottom row: weighted grid graphs.
@@ VII Experiments and Results
This section describes our numerical results of simulations on an Ubuntu 22.04 PC with a 2.50GHz Intel® Core® i5-13490F CPU and 32GB RAM. Our proposed algorithms and the baseline methods are publicly available at Github [^4] https://github.com/reso1/LS-MCPP, with all the benchmark instances and reported numerical results provided as well. In Sec. VII-D, we deployed the proposed algorithms on two Pepper [^5] https://www.aldebaran.com/en/pepper robots in an indoor environment to validate our planning pipeline for MCPP.
@@ VII-A Experiment Setup
Instances: We generate a set of instances to benchmark different methods utilizing the four-way connected 2d undirected grid graphs from previous grid-based MCPP works [15, 21] as well as the 2D pathfinding benchmark [68]. As shown in Fig. 11, the graphs in the top row have unweighted edges with a uniform weight of $1$, whereas the graphs in the bottom row have weighted edges whose weights are randomly generated from $1$ to $3$. Originally, each graph can be contracted to a hypergraph where all its hypervertices are complete. The original graphs are then used to generate 12 groups of 12 mutated MCPP instances (i.e., mutations) of $k$ robots, resulting from a removal ratio $\rho=0,1,...,11$ and a seed $s=0,1,...,11$. Each mutation $(G,I,R)$ with $\rho$ and $s$ is constructed by (1) randomly sampling (with seed $s$) a set of $k$ vertices from $G$ as $R$ and (2) randomly removing (with seed $s$) vertices from $G$ by $x=(\frac{100}{12}\cdot\rho)\%$, meanwhile ensuring the root vertex of any robot and all its neighbors are not removed. The root vertices in $R$ are sampled to ensure they are not cut vertices for constructing well-formed instances [69, 60], that is, a robot can wait indefinitely at its root vertex without blocking any other robot. Note that the graph edge weights of each mutation are also randomly generated using its corresponding seed $s$. For all mutations, we set $C=0.5$ as the $90^{\circ}$ turning cost as described in Sec. VI-D.
Figure 12: Ablation study for single-robot CPP. Each subplot displays the removal ratio $\rho$ (x-axis) versus cost reductions compared to ESTC-UW (y-axis). The cost statistics (mean and variance) are computed across all 12 seeds.
@@
Baselines: We describe all the methods being compared in our experiments. (1) Single-Robot CPP: To verify the effectiveness of our ESTC with the proposed hyperedge weight definition, we use a baseline method, namely ESTC-UW, that computes the circumnavigating coverage path on a spanning tree assuming the edges are unweighted. We also performed an ablation study on the two local optimizations for ESTC: parallel rewiring (PR) in Sec. IV-A and turn reduction (TR) in Sec. IV-C. (2) MCPP: As mentioned in Sec. IV-D, the proposed ESTC algorithm can be integrated into various grid-based MCPP methods, making them immediately applicable to MCPP. To reiterate, VOR, MFC, MIP, and MSTC are the comparing baseline methods with our LS-MCPP. In addition, we also compare LS-MCPP with and without the vertex-wise operators that are new to our previous work [21]  [^6] In the ablation study of our previous work [21], we have shown the effectiveness of grow, deduplicate, and exchange operators as well as the use of Forced Deduplication (Alg. 3) and heuristic sampling for LS-MCPP., namely LS(+VO) and LS(-VO) respectively. Note that all the methods compared use ESTC with TR and PR applied by default. (3) Deconflicted MCPP: Given LS(+VO) solutions to produce the initial time-embedded trajectories, we compare the deconflicting performance using the chaining approach (CHA), the multi-label approach (MLA), and the adaptive approach (ADA) as the low-level planner for PBS.
@@ VII-B Ablation Study
Single-Robot CPP: In Fig. 12, we compare all four ESTC variants with the baseline ESTC-UW. We observe a more critical cost reduction on weighted graphs (upper 6 subplots) than on unweighted graphs (lower 6 subplots). For unweighted graphs, TR results in significant performance improvements since uniformly weighted edges can be aligned in one direction to minimize turning costs. For weighted graphs, TR is much less effective as there are few tie-breaking cases for reducing the turning costs, while PR achieves higher cost reductions as the vertex removal percentage increases. In general, ESTC shows its effectiveness for single-robot CPP, and both PR and TR contribute to substantial path cost reductions.
@@
Figure 13: Performance of LS(+VO) solutions with varying iteration limits. The maximum number of iterations is scaled as $M=b\times\sqrt{|V|/k}$, where different colored lines represent different base scalers $b$.
Figure 14: Ablation study for MCPP. Each subplot displays the removal ratio $\rho$ (x-axis) versus makespan reductions compared to VOR (y-axis). The makespan statistics (mean and variance) are computed across all 12 seeds.
Figure 15: Ablation study for Deconflicted MCPP. Each subplot displays the removal ratio $\rho$ (x-axis) versus makespan increments relative to the input LS(+VO) MCPP solution (left y-axis) and success rates (right y-axis). The makespan statistics (mean and variance) are computed across seeds where CHA, MLA, and ADA all succeed; when certain approach(es) fail in all seeds, statistics are computed across seeds where the remaining approach(es) succeed.
MCPP: We compare the solution quality of MCPP between different methods in Fig. 14. Generally speaking, the mutated instances become more complex as the vertex removal ratio increases. Our LS-MCPP consistently outperforms other methods across all instances, while the performance of the other methods varies, with each showing advantages in different instances. Note that due to the huge memory and runtime requirements of MIP, we only test it on relatively smaller instances (i.e., names with a suffix of small or medium), and its performance is highly restricted by the runtime limit of 1 hour. Specifically for our LS-MCPP, as the vertex removal ratio increases, LS(+VO) performs better with the additional vertex operators compared to LS(-VO). However, when the instances become more complex with a large map and more robots, the solution quality of LS(+VO) gets closer to LS(-VO) due to an inadequate number of local search iterations limited by $M$. To demonstrate that we can further improve the solution quality by saturating LS(+VO) with more iterations, we compare the LS(+VO) solutions with increasing maximum numbers of iterations, as shown in Fig. 13. The solution quality steadily improves as the maximum iteration limit $M$ increases with a linear increase in runtime. In practice, $M$ can be flexibly adjusted according to the available runtime budget, allowing a trade-off between better solution quality and faster planning.
Deconflicted MCPP: We explain the Deconflicted MCPP ablation study results in Fig. 15. CHA demonstrates the highest makespan increments and the lowest success rate due to its inability to backtrack to previous goal states, resulting in incompleteness. This limitation makes CHA susceptible to “dead-ends” when no valid succeeding states are available. Consequently, it reports a higher number of failure nodes during PBS search, which either prevents PBS from finding a valid priority ordering or significantly increases the number of PBS nodes explored (sometimes exceeding the runtime limit in large-scale instances). MLA produces high-quality trajectory sets, with no makespan increases for most of the “easier” instances it can handle within the runtime limit. However, for more complex instances involving larger maps and more robots, the added label dimension across the entire coverage path expands the search space considerably, which significantly increases planning time for each PBS node and results in a lower success rate due to timeouts. Our ADA succeeds in all instances, achieving high solution quality with no makespan increases in most cases. This is due to its ability to restrict the label-augmented search space within a local window, rather than across the entire coverage path as MLA, and only when backtracking to previous states is necessary. In Fig. 16, we also compare the deconflicting results of different MCPP solutions for floor_large and terrain_large using ADA as the low-level planner for PBS. The results demonstrate the better solution quality of the proposed LS-MCPP when used as the input MCPP solution for deconfliction. In contrast, as MFC and MSTC have larger makespans due to the higher overlapping covered regions across the map, the deconflicting procedures on them are more time-consuming. Notably, there is one failure case of deconflicting the MFC solution where the one-hour runtime limit is exceeded. For VOR, as there are no overlapping regions in its solution, PBS immediately returns the initial time-embedded trajectories without any conflict.
Figure 16: Deconflicting different MCPP solutions using ADA as the low-level planner for PBS. The left two subplots show the removal ratio $\rho$ (x-axis) versus makespan (left y-axis) and success rates (right y-axis). The makespan statistics (mean and variance) are computed across seeds where ADA succeeds for all input MCPP solutions.
VII-C Planning Runtime Comparison
We report the planning runtime of all the MCPP methods and the Deconflicted MCPP methods across all the 12 groups of instances in Fig. 11. We show the planning runtime comparison of methods for MCPP in Tab. I, where our LS-MCPP requires more time to explore the solution neighborhood. In particular, adding vertex-wise operators only increases the runtime of LS-MCPP by a little if we compare LS(-VO) and LS(+VO). Nevertheless, it produces high-quality MCPP solutions within around 13 minutes even for large-scale instances such as NewYork1 and Shanghai2, which is acceptable for offline planning in real-world robotics applications.
@@ VII-D Physical Robot Experiments
We constructed an MCPP instance of an indoor environment with a Vicon [^7] https://www.vicon.com motion capture system for localization, as shown in Fig. 17. Our pipeline computes the conflict-free set of trajectories for the robots by first finding an MCPP solution via LS-MCPP and then deconflicting the MCPP solution via PBS and ADA. After offline planning, the trajectories are sent to the robots to execute the coverage task via a centralized scheduler implemented in ROS [^8] https://www.ros.org. Fig. 17 demonstrates the execution procedure of the resulting trajectories using our planning pipeline. The detailed setup is described as follows.
@@
Multi-Robot System: We employ two Pepper robots, named Cayenne and Cumin, each with a square bounding box of 0.55 meters. The 2D grid graph represents the indoor environment as a $10\times 6$ map with each grid unit sized at 0.65 meters to accommodate their bounding boxes. The initial root vertices of Cayenne and Cumin are the bottom-left and top-right cells, respectively. Both Peppers can have omnidirectional movement driven by velocity commands $u=(v_{x},v_{y},\omega)$. The linear velocities $v_{x}$ and $v_{y}$ are capped at 0.15 meters per second, while the angular velocity $\omega$ is restricted to 0.5 radians per second. To determine the non-uniform traversal costs (i.e., the edge weights) for Peppers, we recorded the maximum and minimum time costs to traverse a single grid unit by performing random move actions across the graph. We then sampled weights randomly within this range and assigned them to the edges. Although Peppers can move omnidirectionally, we assume they always face the moving directions by integrating turn actions as described in Sec. VI-D. We set $C$ as the maximum action time of 90 rotations for Peppers. In the final constructed graph, the time costs for move actions range from 7.2 to 10.8 seconds, while the time cost for a 90 turn is 6.0 seconds.
Figure 17: Demonstration of Cayenne and Cumin executing the coverage task cooperatively and without conflicts. Four frames from the execution are shown, with robot trajectories and the grid graph displayed on the left. Static obstacles and the trajectories of Cayenne and Cumin are highlighted in red, blue, and yellow, respectively. The entire planning pipeline was completed in 0.1 seconds, and the execution of the coverage task took approximately 6 minutes.
@@
Centralized Scheduler: Given the set of conflict-free trajectories resulting from our planning pipeline, we generate an action sequence consisting of move, turn, and wait actions between consecutive states for each Pepper. A centralized scheduler is employed to sequentially send actions to the corresponding Pepper, maintaining a positional tolerance of 0.01 meters and a rotational tolerance of 0.02 radians. The velocity commands for each action are generated using a proportional controller operating at 20 Hz to align with the nominal action time cost. In practice, the move and turn action time costs are not precise for the Peppers, as we assume instantaneous velocity changes and do not account for their real-time dynamics or potential perturbations. This often leads to an overestimation of the actual action time costs, as the time for acceleration and deceleration is ignored, though underestimations can also occur occasionally. Therefore, we design the centralized scheduler to synchronize the discrepancies between the nominal time costs $t$ and the actual time costs $t^{\prime}$ for the actions. Upon completing each action, each Pepper sends the time difference $\Delta t=t-t^{\prime}$ back to the scheduler. If $\Delta t<0$, the scheduler signals the Pepper to wait for a period of $-\Delta t$ to match the nominal time cost; otherwise, the scheduler instructs the other Pepper to add $\Delta t$ to its waiting period after completing its current action.
@@
Our centralized scheduler currently uses a basic approach to synchronize time differences between nominal and actual action time costs, which may increase the task completion time. More advanced techniques from MAPF execution research, such as constructing an action dependency graph [70], could address this issue. However, exploring these advanced methods is beyond the scope of this paper.